In April 1974, Steve Prefontaine completed a 10 km race in a time of 27 min, 43.6 s. Suppose "Pre" was at the 8.27 km mark at a time of 25.0 min. If he accelerated for 60 s and then maintained his increased speed for the remainder of the race, calculate his acceleration over the 60 s interval. Assume his instantaneous speed at the 8.27 km mark was the same as his overall average speed up to that time.

Please explain, I have gotten many answers and its been wrong every time.
Answer #1

Total distance covered in the race d 10 km

Total time taken to complete the race t -27 min 43.6 s

Time taken to complete distance dı 8.27 kım is = 25 min

Average speed during this time is 8.27* 10 5.513m/s t1 25 60 u -

Remaining distance is d2 + d3 = 10-8.27-1.73 km-1730 m -----(1)

Distance d_2 is completed with an acceleration a in a time t_2=60,s

d_2=ut_2+rac{1}{2}at_2^2 ----(2)

Cl t2 ---(3)

Distance d_3 is completed with average speed v in a time t_3=27min,43.6,s-25min-60,s=1min43.6,s=103.6,s

v t3 (l3 ---(4)

using (2) and (4) in (1)

ut_2+rac{1}{2}at_2^2+vt_3=1730 ------(5)

using equations (3) and (4)

ut_2+rac{1}{2}at_2^2+(u+at_2)t_3=1730

a=rac{2*1730-2ut_2-2ut_3}{t_2^2+2t_2t_3}

substituting the values of u, t2,t3 in the above equation,

a=rac{2*1730-2*5.513*60-2*5.513*103.6}{60^2+2*60*103.6}

a0.1033m/s